# Cartesian Forest Matching

### v1

@@ 8       Update $LN_{x}$
Algorithm 1 solves the CFM problem.
Proof
In Algorithm 1, that solves the CFM problem, $LN_{x}$ is a linear representations of the Cartesian Forest of a sequence $x$. It can indifferently be its Parent-Distance table or its Skipped-number representation. Note that this algorithm already appears in [1].
In [1] the authors show that this algorithm (originally applied to solve the Cartesian Tree Matching problem) has a $\mathcal{O}(mn)$ worst-case time-complexity and a $\mathcal{O}(m)$ space-complexity. It is also proved that the average-case time-complexity is $\mathcal{O}(n)$ in several random models.
The same results hold for Cartesian Forest Matching (even if some average-case results require an adaptation of the proofs, the main ideas remain).
@@ 4 Approximate Cartesian Forest matching
The above approach for exact CFM can be extended to approximate CFM with one difference in a similar way as approximate CTM with one difference is done in [1].
Given a sequence $x$ of length $m$ and $i$ a position $1\leq i<m$, let $y$ be the sequence defined by swapping the elements in positions $i$ and $i+1$ in $x$. In [1], it is shown that there are at most $3$ mismatches between the Skipped-number representation of the Cartesian Tree of $x$ and the Skipped-number representation of the Cartesian Tree of $y$. The same result holds for the Skipped-number representation of the Cartesian Forest of $x$ and the Skipped-number representation of the Cartesian Forest of $y$ considering that two elements match if their absolute values are equal. Thus, based on the results in [1], the CFM with one swap of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
Approximate CTM with one mismatch, one insertion or one deletion is solved by comparing the Parent-Distance representations from left to right and the Parent-Distance representations from right to left of the Cartesian Trees of $x$ and $y$. The Parent-Distance representation from right to left of a Cartesian Forest can be defined in a similar way as for a Cartesian Tree. And then, approximate CFM with one mismatch, one insertion or one deletion can be solved in a similar way than approximate CTM with one mismatch, one insertion or one deletion. Thus, based on the results in [1], the CFM with one mismatch, one insertion or one deletion of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
@@ 5.1 Recursive Definition
According to Definition 2, a Cartesian Forest $F$ can be either empty (denoted by $\emptyset$) or contains at least one node. This node has a left sub-forest and a right sub-forest, but it can also have siblings. A sibling $S$ is a particular Cartesian Forest that cannot have a left sub-forest. Therefore, we obtain the following recursive decomposition:
@@ 5.2 Generating Function
Let $F(z)=\sum_{n\geq 0}f_{n}z^{n}$ be the generating function of Cartesian Forests, where $f_{n}$ counts the number of Cartesian Forests with $n$ nodes and $S(z)$ be the generating function of siblings. From the previous recursive decomposition we have:
@@
from which we obtain that
$S(z)=\frac{1}{1-z\cdot F(z)}\text{ and }F(z)=1+\frac{z\cdot F^{2}(z)}{1-z\cdot F(z)}$
It can be shown from here that $F(z)=\frac{1+z+\sqrt{1-6z+z^{2}}}{4z}$, whose associated coefficients are known to be Schröder–Hipparchus numbers, also called super-Catalan numbers (Sequence $A001003$ on OEIS) enumerated by the following formula:
$f_{n}=\sum_{i=1}^{n}\frac{1}{n}\binom{n}{i}\binom{n}{i-1}2^{i-1}$
which, amongst other things, counts the number of ways of inserting parentheses into a sequence of $n+1$ symbols, where each pair of parentheses surrounds at least two symbols or parenthesized groups, and without any parentheses surrounding the entire sequence. In [2], the authors show that:
@@
In the following subsections, we describe bijections between Cartesian Forests and classical combinatorial objects counted by the Schröder-Hipparchus Number. Figure 5 shows the one to one correspondence for $n\in\{1,2,3\}$ between Cartesian Trees, Schröder Trees and Parentheses Words.
@@ 5.3 A bijection with Schröder Trees
A Schröder Tree is a planar tree whose internal nodes have two or more subtrees.
Schröder Trees with $n+1$ leaves are counted by $f_{n}$.
Given a Cartesian Forest $F$ with $k$ roots $(r_{1},\ldots,r_{k})$, we denote $\operatorname{left}(r_{1})$ the left sub-forest of the root $r_{1}$. For $i\in\{1,\ldots,k\}$, we denote $\operatorname{right}(r_{i})$ the right sub-forest of $r_{i}$.
From a Cartesian Forest with $n\geq 1$ nodes, Algorithm 2 gives a recursive approach to building a Schröder Tree with $n+1$ leaves .
Output : A Schröder Tree $ST$ with $n+1$ leaves
@@ 8      $ST\leftarrow$ Add subtrees $(c_{1},\ldots,c_{k+1})$ to $ST$;
10return $ST$;
Algorithm 2 CFtoST$(F)$
@@
Output : A sequence $w$ with parentheses
@@ 8      $w\leftarrow\bm{(}w_{1}\cdot w_{2}\cdots w_{k+1}\bm{)}$;
Figure 4: Both algorithms take a Cartesian Forest as input and return the corresponding Schröder Tree (on the left) or Parentheses Word (on the right). As one can see, both methods are very similar.
@@ 5.4 A bijection with Parentheses Words
A Parentheses Word $w$ is a word: over the alphabet {$\{\bm{(}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \square\leavevmode\nobreak\ ,\leavevmode\nobreak\ \bm{)}\}$ such that
  • either $w=\square$
  • or $w=\bm{(}w_{1}\cdots w_{k}\bm{)}$ where $k\geq 2$ and each $w_{i}$ is a parentheses word.
Do note that, in this definition, unlike the more commonly found definition of these words, we allow parentheses to surround the entire sequence. We notably do so in order to simplify Algorithm 3 and to make the bijection more apparent to the reader. But since every parentheses word that is not . is contained between a common parentheses, it is not necessary to represent it. Hence, in Figure 5, we omit to draw those parentheses in order to match classical representations (we write $\bm{(}\square\square\bm{)}\square$ instead of $\bm{(}\bm{(}\square\square\bm{)}\square\bm{)}$ for instance). Informally, one may consider the symbols of the word as separators between the nodes and each group of parentheses as a (sub-)forest.
Using the same arguments as for Schröder Tree, Algorithm 3 is an injective function.
@@
Algorithms 2 and 3 are bijective functions that map a Cartesian Forest with $n$ nodes to a Schröder Tree with $n+1$ leaves (Algorithm 2) or to a Parentheses Word with $n+1$ symbols (Algorithm 3). They both have a $\Theta(n)$ time-complexity and a $\Theta(n)$ worst-case space complexity.
Figure 5: Correspondence between Cartesian Forests with $n$ nodes, Schröder Trees with $n+1$ leaves and sequences of length $n+1$ with parentheses when $n\in\{1,2,3\}$.
@@ 6 Cartesian Forest Signature and Cartesian Forest Matching using a filter
In [4], the authors propose a signature, or perfect hash, of a Cartesian Tree, based on the Skipped-number representation. Given a Cartesian Tree with $n$ nodes, its signature is an integer with at most $2n$ bits. In this section, we extend this notion of signature to Cartesian Forests, obtaining a signature with at most $3n$ bits.
@@
Given a Cartesian Forest $F$ and its Skipped-number representation $SN_{F}$ (see Definition 6), its signature is defined in the following way: given a position $i\in\{1,\ldots,m\}$ in the Skipped-number representation, each value $SN_{F}[i]$ is encoded by a sequence of bits, that are concatenated to obtain an integer.
  • The first bits concern the sign of $SN_{F}[i]$. If $SN_{F}[i]=0$, it is equal to 00. The case $SN_{F}[i]<0$ is encoded by $10$ and $SN_{F}[i]>0$ is encoded by $11$.
  • If $SN_{F}[i]\neq 0$, the following bits are a unary encoding of $|SN_{F}[i]|$: $|SN_{F}[i]|-1$ bits equal to $1$ followed by a bit equal to 00.
Since the total number of skipped nodes cannot exceed $m$, a signature contains at most $3m$ bits.
This representation could be used efficiently as a perfect hash for small patterns, to obtain an efficient algorithm to solve the Cartesian Forest Matching problem: one only needs to update the signature as one would update the Skipped-number representation of the chunk of text compared to the pattern, using bitwise operations. But it becomes inefficient when $3m$ exceeds the size of a register, as one cannot use those operations efficiently anymore.
Therefore, in order to accelerate the Cartesian Forest Matching in a Rabin-Karp fashion we now introduce a filtering method.
@@
Given a sequence $x$ of length $m$ and its linear representation $LN_{x}$, its $\tau$-filter $\mathcal{F}il_{x}$ is a sequence of $\tau$ bits such that $\mathcal{F}il_{x}[i]=0$ if $LN_{x}[m-\tau+i]=0$ and $\mathcal{F}il_{x}[i]=1$ otherwise.
Algorithm 1 can be adapted: line $5$ is only tested if $\mathcal{F}il_{x}$ is equal to $\mathcal{F}il_{p}$. $\tau$ can be adapted to match the size of a classical register, that is $32$, $64$ or $128$ bits. Therefore, comparing two filters can be made in constant time. The update of the filter can also be made in constant time by tracking the positions in $LN_{x}$ that have been updated (see [1] for more details on the update function of the linear representations). In the following Section, we only implemented the filter for the Skipped-number representation.
@@ 7 Experiments
In this Section, we implemented a Cartesian $\tau$-filter using $\tau=64$. The first set of experiments consider the uniform distribution over text and patterns of respective length $n$ and $m$, over a $k$ letter alphabet. Figure 6 sums up the details and observations.
Figure 6: In this experiment, we randomly generated 10 0001000010\,00010 000 patterns of size $m$ and texts of size $1000$. From left to right, the alpĥabet is respectively equal to $2$, $4$ and finally $m$. As one, can see, the Parent-Distance version is the slowest, then comes the one using the Skipped-number representation. The method using a filter is the faster one, but only with a slight margin. The average cost slightly increases with $k$, which is probably due to the cost of the update function, since the average Parent-Distance increases with the size of the alphabet. On the contrary, the standard deviation seems to decreases, since we obtain smoother curves.
Efficient even with (not too) low entropy
Figure 7: Both figures represent the same experiment: the one on the left is the full one and the one on the right is a zoom on the results when the entropy is low (between $0.01$ and $0.3$). In this experiment, for each value of the collision entropy, 10 0001000010\,00010 000 random texts of length $1000$ and patterns of length $100$ were generated and the three algorithms were applied to them. The added value of using the filter method, compared to the simple Skipped-number representation version, is more important when the entropy is low, whereas the difference in efficiency between the Parent-Distance version and the Skipped-number representation version drops.
@@
The following experiment uses the random generator from [3]. It generates sequences uniformly amongst those of a fixed length over a fixed alphabet and a fixed Collision Entropy (that is Rényi Entropy with $\alpha=2$). The Collision Entropy is a function of the probability for two random variables (a letter in the pattern $p$ and the text $t$) to be equal. The lower the entropy is, the higher average complexity of the algorithms. As a matter of fact, if the entropy reaches 00, then both the text and the pattern contains the repetition of a unique symbol, which corresponds to the worst-case complexity of Algorithm 1. Though, as one can see in Figure 7, the average cost of the algorithms quickly drops and stabilizes.

### v2

@@ 4 Exact Cartesian Forest Matching
Algorithm 1 solves the CFM problem in $\mathcal{O}(n)$ space and $\mathcal{O}(mn)$ time in the worst-case and $\mathcal{O}(n)$ time on average.
Proposition 2
Using a Cartesian Forest border table, the CFM problem can be solved in $\mathcal{O}(n)$ time and space by adapting the KMP approach used in [19].
Note that, in order to prove both propositions, it is sufficient to prove that two sequences share the same Cartesian Forest if and only if they share the same linear representation. It can done by adapting the proof from [19]111see Appendix A.1.
@@ 5 Approximate Cartesian Forest Matching
The above approach for exact CFM can be extended to approximate CFM with one difference in a similar way as approximate CTM with one difference is done in [3]. Given a sequence $x$ of length $m$ and $i$ a position $1\leq i<m$, let $y$ be the sequence defined by swapping the elements in positions $i$ and $i+1$ in $x$. In [3], it is shown that there are at most $3$ mismatches between the Skipped-number representation of the Cartesian Tree of $x$ and the Skipped-number representation of the Cartesian Tree of $y$. The same result holds for the Forest Skipped-Number representation of the Cartesian Forest of $x$ and the Forest Skipped-Number representation of the Cartesian Forest of $y$ considering that two elements match if their absolute values are equal. Thus, based on the results in [3], the CFM with one swap of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
Approximate CTM with one mismatch, one insertion or one deletion is solved by comparing the Parent-Distance representations from left to right and the Parent-Distance representations from right to left of the Cartesian Trees of $x$ and $y$. The Forest Parent-Distance representation from right to left of a Cartesian Forest can be defined in a similar way as for a Cartesian Tree. And then, approximate CFM with one mismatch, one insertion or one deletion can be solved in a similar way than approximate CTM with one mismatch, one insertion or one deletion. Thus, based on the results in [3], the CFM with one mismatch, one insertion or one deletion of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
@@ 6.1 Recursive Definition
Definition 10(Cartesian Forest $F$)
A Cartesian Forest $F$ can be:
  • empty,
  • a sequence of $k$ trees rooted in $(r_{1},\ldots,r_{k})$, with $k\geq 1$, such that $r_{1}$ has a left Cartesian sub-Forest and a right Cartesian sub-Forest, and $r_{i}$ has only a right Cartesian sub-Forest for all $i\in\{2,\ldots,k\}$ (the left sub-Forest is necessarily empty).
Figure 4: On the left is a valid Cartesian Forest. In the middle, it is not a valid Cartesian Forest because the second tree in the right sub-forest of $r^{\prime}_{1}$ has a left sub-forest. On the right, the forest is not Cartesian: the second tree has a left sub-forest.
@@ 6.2 Generating Function
Let $F(z)=\sum_{n\geq 0}f_{n}z^{n}$ be the generating function of Cartesian Forests, where $f_{n}$ counts the number of Cartesian Forests with $n$ nodes, and $S(z)$ be the generating function of siblings. The translation between the recursive decomposition is made using the Symbolic Method [8]:
@@
from which we obtain that $S(z)=\frac{1}{1-z\cdot F(z)}$ and $F(z)=1+\frac{z\cdot F^{2}(z)}{1-z\cdot F(z)}$.
It can be shown from here that $F(z)=\frac{1+z+\sqrt{1-6z+z^{2}}}{4z}$, whose associated coefficients are known to be Schröder–Hipparchus numbers, also called super-Catalan numbers (Sequence $A001003$ on OEIS) enumerated by the following formula: $f_{n}=\sum_{i=1}^{n}\frac{1}{n}\binom{n}{i}\binom{n}{i-1}2^{i-1}$, which, amongst other things, counts the number of ways of inserting parentheses into a sequence of $n+1$ symbols, where each pair of parentheses surrounds at least two symbols or parenthesized groups, and without any parentheses surrounding the entire sequence. In [4], the authors show that:
@@
In the following subsections, we describe bijections between Cartesian Forests and classical combinatorial objects counted by the Schröder-Hipparchus numbers. We recall that since those objects all share the same generating function, there exists a bijection between those sets [^2] Figure 8 (Appendix) shows the one to one correspondence for $n\in\{1,2,3\}$ between Cartesian Trees, Schröder Trees and Parentheses Words. Any injection from one set to the other is a bijection.
@@ 6.3 A bijection with Schröder Trees
A Schröder Tree is a tree whose internal nodes have two or more subtrees.
Schröder Trees with $n+1$ leaves are counted by $f_{n}$. Given a Cartesian Forest $F$ with $k$ roots $(r_{1},\ldots,r_{k})$, we denote $\operatorname{left}(r_{1})$ the left sub-forest of the root $r_{1}$. For $i\in\{1,\ldots,k\}$, we denote $\operatorname{right}(r_{i})$ the right sub-forest of $r_{i}$. From a Cartesian Forest with $n\geq 1$ nodes, Algorithm 2 gives a recursive approach to building a Schröder Tree with $n+1$ leaves.
Input : A Cartesian Forest $F$ w. $n$ nodes and $k$ roots, s.t. $F=(r_{1},\ldots,r_{k})$
Output : A Schröder Treewith $n+1$ leaves
@@ 8 $ST\leftarrow$ Add subtrees $(c_{1},\ldots,c_{k+1})$ to $ST$;
10return STST;
Algorithm 2 CFtoST(F)(F)
@@
Output : A sequence with parentheses and n+1n+1\ $\square$
@@ 8 $w\leftarrow\bm{(}w_{1}\cdot w_{2}\cdots w_{k+1}\bm{)}$;
Figure 5: Algorithms that take a Cartesian Forest as input and return the corresponding Schröder Tree (algorithm on the left) or Parentheses Word (algorithm on the right). As one can see, both methods are very similar. In both cases, two different inputs imply two different outputs. Hence both functions are injections. Since there is the same number of objects of the same size in all three cases, the functions are bijections.
@@ 6.4 A bijection with Parentheses Words
[…]
Do note that, in this definition, unlike the more commonly found definition of these words, we allow parentheses to surround the entire sequence. We notably do so in order to simplify Algorithm 3 and to make the bijection more apparent to the reader. But since every parentheses word that is not $\square$ is contained between a pair of parentheses, it is not necessary to represent it. Informally, one may consider the symbols of the word as separators between the nodes and each group of parentheses as a (sub-)forest. Using the same arguments as for Schröder Tree, Algorithm 3 is an injective function.
@@
Algorithms 2 and 3 are bijective functions that map a Cartesian Forest with $n$ nodes to a Schröder Tree with $n+1$ leaves (Algorithm 2), and to a Parentheses Word with $n+1$ symbols (Algorithm 3). They both have a $\Theta(n)$ time-complexity and a $\Theta(n)$ worst-case space complexity.
@@ 7 Cartesian Forest Signature and Cartesian Forest Matching using a filter
In [6], the authors propose a signature (or perfect hash) of a Cartesian Tree, based on the Skipped-number representation. Given a Cartesian Tree with $n$ nodes, its signature is an integer with at most $2n$ bits. In this section, we extend this notion of signature to Cartesian Forests, obtaining a signature with at most $3n$ bits.
@@
Given the Forest Skipped-Number representation $\operatorname{F-SN}_{x}$ of a sequence $x$ (see Definition 8), its signature is defined in the following way: given a position $i\in\{1,\ldots,n\}$ in the Forest Skipped-Number representation, each value $\operatorname{F-SN}_{x}[i]$ is encoded by a sequence of bits, that are concatenated to obtain an integer.
  • The first bits concern the sign of $\operatorname{F-SN}_{x}[i]$: If $\operatorname{F-SN}_{x}[i]=0$, it is equal to $\tt 0$}; The case $\operatorname{F-SN}_{x}[i]<0$ is encoded by $\tt 10$}, and the case $\operatorname{F-SN}_{x}[i]>0$ is encoded by $\tt 11$}.
  • If $\operatorname{F-SN}_{x}[i]\neq 0$, the following bits are a unary encoding of $|\operatorname{F-SN}_{x}[i]|$: $|\operatorname{F-SN}_{x}[i]|-1$ bits equal to $\tt 1$} followed by a bit equals to $\tt 0$}.
Since the total number of skipped nodes cannot exceed $n$, a signature contains at most $3n$ bits. This representation could be used efficiently as a perfect hash for small patterns, to obtain an efficient algorithm to solve the Cartesian Forest Matching problem: one only needs to update the signature as one would update the Forest Skipped-Number representation of the chunk of text compared to the pattern, using bitwise operations. But it becomes inefficient when $3n$ exceeds the size of a register, as one cannot use those operations efficiently anymore. Therefore, in order to accelerate the Cartesian Forest Matching in a Rabin-Karp fashion we now introduce a filtering method.
@@
Given a sequence $x$ of length $n$ and its linear representation $\operatorname{LN}_{x}$, its $\tau$-Filter, denoted by $\mathcal{F}il_{x}$, is a sequence of $\tau$ bits such that $\mathcal{F}il_{x}[i]=0$ if $\operatorname{LN}_{x}[n-\tau+i]=0$ and $\mathcal{F}il_{x}[i]=1$ otherwise.
Algorithm 1 can be adapted: line $5$ is only tested if $\mathcal{F}il_{x}$ is equal to $\mathcal{F}il_{p}$. $\tau$ can be adapted to match the size of a classical register, that is $32$, $64$ or $128$ bits. Therefore, comparing two filters can be made in constant time. The update of the filter can also be made in constant time by tracking the positions in $\operatorname{LN}_{x}$ that have been updated (see [3] for more details on the update function of the linear representations). In the following Section, we only implemented the filter for the Forest Skipped-Number representation.
@@ 8 Experiments
In this Section, we implemented a Cartesian Forest $\tau$-Filter using $\tau=64$. The first set of experiments considers the uniform distribution over text and patterns of respective length $n$ and $m$, over a $k$ letter alphabet. Figure 6 sums up the details and observations. As one can see, the Forest Parent-Distance version is the slowest (PD meta), then comes the one using the Forest Skipped-Number representation (SN meta). The method using a filter (SN hash) is even faster, but only with a slight margin. Finally, the KMP version [19] is the fastest one (PD KMP). Note that we tried to implement the Skip-Search algorithm [21] without the SIMD instructions, but the experimental results were disappointing. The average cost slightly increases with $k$, which is probably due to the cost of the update function, since the average Forest Parent-Distance increases with the size of the alphabet. Note that, whatever the linear representation is, the number of comparisons at Line $5$ in Algorithm 1 is the same, since they encode the same prefix. Therefore, it is likely that the extra cost of the Forest Parent-Distance representation is due to the fact that both requires to use a stack in their implementations but more values are pushed in the stack, in the Forest Parent-Distance case.
Figure 6: We randomly generated 10 00010\,000 patterns of size $m$ and texts of size 1 0001\,000. The $x$-axis represents $m$ and the $y$-axis is the average time taken by each algorithm in microseconds. From left to right, the alpĥabet size is respectively equal to 22, 44 and $m$.
@@
The experiment in Figure 7 uses the random generator from [5]. It shows the algorithm remains efficient even with (not too) low entropy. It generates sequences uniformly amongst those of a fixed length over a fixed alphabet and a fixed Collision Entropy (that is Rényi Entropy with $\alpha=2$). The Collision Entropy is a function of the probability for two random variables (a letter in the pattern $p$ and the text $t$) to be equal. The lower the entropy is, the higher average complexity of the algorithms. As a matter of fact, if the entropy reaches 0, then both the text and the pattern contains the repetition of a unique symbol, which corresponds to the worst-case complexity of Algorithm 1. Though, as one can see in Figure 7, the average cost of the algorithms quickly drops and stabilizes. The KMP version is the most efficient (this is in line with the results of [2] showing that the KMP algorithm is efficient for order preserving matching) and is (unsurprisingly) very stable, even when the entropy is low.
Figure 7: For each value of the collision entropy ($x$-axis), 10 00010\,000 random texts of length 1 0001\,000 and patterns of length $100$ were generated, and the four algorithms were applied to them. On the left $m=10$, and on the right $m=100$. The $y$-axis is the average time in microseconds. The added value of using the filter method, compared to the simple Forest Skipped-Number version, is more important when the entropy is low, whereas the difference in efficiency between the Forest Parent-Distance version and the Forest Skipped-Number version drops.
Appendix 0.A Appendix
0.A.1 On the Linear Representations of Cartesian Forests
We show that two sequences $x$ and $y$ share the same Cartesian Forest if and only if they share the same Forest Parent-Distance representation.
Proof
If the sequences have different lengths, then they have different Cartesian Forests and different Forest Parent-Distance representations and it holds. We need only consider the case when $x$ and $y$ have the same length $m$. We use a proof by induction on $m$.
For $m=0$, both Cartesian Forests and their representations are empty and trivially match. For $m=1$, both Cartesian Forests have only one root and are the same, and their Forest Parent-Distance representation are both equal to 0. The property holds for $m=1$. Let us now assume that the property holds for a certain $m=k$ and move onto $m=k+1$.
[…]
  • If positions $k+1$ have a parent in both Cartesian Forests, let $\alpha$ (resp. $\beta$) be the parent of $k+1$ in $F(x[1\ldots k+1])$ (resp. $F(y[1\ldots k+1])$). Since both sequences share the same Cartesian Forest, we must have $\alpha=\beta$. We have $\operatorname{F-PD}_{x}[1\ldots k]=\operatorname{F-PD}_{y}[1\ldots k]$ by induction hypothesis. If $\alpha$ and $\beta$ are parents then $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]=k+1-\alpha=k+1-\beta$. Therefore $\operatorname{F-PD}_{x}[1\ldots k+1]=\operatorname{F-PD}_{y}[1\ldots k+1]$.
  • If positions $k+1$ have siblings in both Cartesian Forests, let $\alpha$ (resp. $\beta$) be the sibling of $k+1$ in $F(x[1\ldots k+1])$ (resp. $F(y[1\ldots k+1])$). The same reasoning as before applies, except $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]=-(k+1-\alpha)=-(k+1-\beta)$. Therefore $\operatorname{F-PD}_{x}[1\ldots k+1]=\operatorname{F-PD}_{y}[1\ldots k+1]$.
  • If positions $k+1$ possess neither parents or siblings, then we have $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]=0$ and thus $\operatorname{F-PD}_{x}[1\ldots k+1]=\operatorname{F-PD}_{y}[1\ldots k+1]$.
[…]
  • If $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]\neq 0$, let $\alpha=k+1-|\operatorname{F-PD}_{x}[k+1]|$.
    • If $\operatorname{F-PD}_{x}[k+1]<0$, then $\alpha$ is the sibling of $k+1$ in both $F(x[1\ldots k+1])$ and $F(y[1\ldots k+1])$.
    • Otherwise, if $\operatorname{F-PD}_{x}[k+1]>0$, that means $\alpha$ is the parent of $k+1$ in $F(x[1\ldots k+1])$ and $F(y[1\ldots k+1])$. Also the right sub-Forest of $\alpha$ becomes the left sub-Forest of $k+1$.
  • If $\operatorname{F-PD}_{x}[k+1]=0$, $k+1$ is the root of $F(x[1\ldots k+1])$ and $F(x[1\ldots k])$ becomes the left sub-Forest of $x[k+1]$. We derive $F(y[1\ldots k+1])$ from $F(y[1\ldots k])$ in the same way.
In all three cases, we can conclude that $F(x[1\ldots k+1])=F(y[1\ldots k+1])$.
Hence, there is a one-to-one mapping between Cartesian Forests and Forest Parent-Distance representations.
A similar proof using the same arguments can be used to prove a one-to-one mapping between Cartesian Forests and the Forest Skipped-Number representation.
0.A.2 On the Symbolic Method
The Symbolic Method was introduced in [8]. It is a method to directly translate a recursive definition of a combinatorial class into a generation function. Recall that a combinatorial class is a set of objects associated with a size function, such that for all $n\geq 0$, the subset of objects of size $n$ is finite. This section only contains the information necessary to understand how we obtained the generating function $F(z)$. Therefore, we focus on unlabelled objects and simple operations. A reader willing to learn more should read [8].
Let $\mathcal{C}$ be a combinatorial class and $\mathcal{C}_{n}$ be the set of objects of $\mathcal{C}$ of size $n$. We have $\mathcal{C}=\bigcup_{n\geq 0}\mathcal{C}_{n}$ The generating function $C(z)$ of $\mathcal{C}$ is defined as a power series
$C(z)=\sum_{n\geq 0}c_{n}z^{n},$
where $c_{n}=|\mathcal{C}_{n}|$.
The symbolic method is a dictionary that translates operations on sets into operations onto their associated generating functions.
The Null size element.
Assume $\mathcal{C}=\mathcal{C}_{0}$, that is it only contains an object of size 0. Then $C(z)=1$.
The size 1 element.
Assume $\mathcal{C}=\mathcal{C}_{1}$, that is it only contains an object of size $1$. In our case, this would be a tree containing a unique node. Then $C(z)=z$.
Disjoint Union.
Assume $\mathcal{C}=D\cup E$, where $D$ and $E$ are two combinatorial classes. Then $C(z)=D(z)+E(z)$.
Cartesian Product
Assume $\mathcal{C}=D\times E$, where $D$ and $E$ are two combinatorial classes. Then $C(z)=D(z)\cdot E(z)$.
We recall the recursive decomposition that we obtained for Cartesian Forest:
[…]
The first combinatorial class $F$ is defined as:
  • either an empty forest (marked with $\emptyset$ whose associated generating function is $1$).
  • or a Cartesian product between a sequence of siblings, whose generating function will be denoted $S(z)$ and a tree, that is decomposed into three parts:
    • a root, that is a tree of size $1$, whose generating function is $z$
    • a left and a right sub-forest, that is to say two Cartesian Forests.
    The generating function of this tree is therefore equal to $F(z)\cdot z\cdot F(z)$.
Using the Symbolic Method we therefore obtain that $F(z)=1+z\cdot F^{2}(z)\cdot S(z)$. Using the same ideas, we obtain that
$S(z)=z\cdot F(z)\cdot S(z)+1$
$\implies S(z)-z\cdot F(z)\cdot S(z)=1$
$\implies S(z)=\frac{1}{1-z\cdot F(z)}$
And therefore
$F(z)=1+\frac{z\cdot F^{2}(z)}{1-z\cdot F(z)}$
$\implies F(z)\cdot(1-z\cdot F(z))=1-z\cdot F(z)+z\cdot F^{2}(z)$
$\implies F(z)-z\cdot F^{2}(z)=1-z\cdot F(z)+z\cdot F^{2}(z)$
$\implies 2z\cdot F^{2}(z)-(1+z)\cdot F(z)+1=0$
Solving this equation we obtain $\Delta=b^{2}-4ac=(1+z)^{2}-8z=1-6z+z^{2}$ and
$F(z)=\frac{-b+\sqrt{\Delta}}{2a}=\frac{1+z+\sqrt{1-6z+z^{2}}}{4z}$
Note here that we do not consider $\frac{-b-\sqrt{\Delta}}{2a}$ because we know $F(z)$ must have positive coefficients.
0.A.3 Bijection with Schröder Trees and Parentheses Words
[…]
Figure 8: Correspondence between Cartesian Forests with $n$ nodes, Schröder Trees with $n+1$ leaves and sequences of length $n+1$ with parentheses when $n\in\{1,2,3\}$ (respect. top, middle, bottom). $\square$ symbols represent leaves and $\circ$ symbols represent internal nodes.